OpenStudy (sophadof):

Write a quadratic function in standard form having zeros of -5 and 1.

OpenStudy (math&ing001):

Use this f(x)=(x-a)(x-b) where a and b are your zeros.

OpenStudy (sophadof):

I don't get what that equation is though.. @math&ing001

OpenStudy (math&ing001):

Example suppose your zeros are 9 and -7 f(x) = (x-9)(x-(-7)) = (x-9)(x+7) = x^2+7x-9x-63 = x^2 - 2x - 63 so f(x)= x^2 - 2x - 63 is the quadratic function in standard form

OpenStudy (sophadof):

oh ok!!

OpenStudy (sophadof):

would this one be x^2+4-5?

OpenStudy (sophadof):

@math&ing001

OpenStudy (math&ing001):

Yep ! Don't forget the x near 4.

OpenStudy (sophadof):

oh thank you!!! <3